# Floquet Schrieffer-Wolff transform based on Sylvester equations

### v2

@@ VI Acknowledgement
The authors would like to thank André Eckardt, Thomas Pohl, Francesco Petiziol, Alexander Schnell, and Li You for their helpful discussions on the spurious Floquet-mediated interactions and the relation with van Vleck theory. DJ and FS acknowledge support from the Cluster of Excellence ’Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994.
@@ A.1 Link to the retarded Green function
where $\theta_{t}$ is the step function at $t=0$, $\big{(}\hat{H}_{1}^{(1)}\big{)}^{H}_{t}=e^{i\hat{H}^{(0)}t}\hat{H}_{1}^{(1)}e^{-i\hat{H}^{(0)}t}$ represents the Heisenberg operator of $\hat{H}_{1}^{(1)}$. Hence, $\langle[\hat{f}^{(1)}_{1},\hat{A}]\rangle$ is a frequency-domain retarded Green function of the undriven system $\hat{H}^{(0)}$. This direct link between $\hat{f}^{(n)}_{j}$ and the retarded Green function suggests that we can obtain the linear and non-linear responses of the system from the solutions of the Sylvester equations   [^5] For example, in a monochromatically driven system described by $\hat{H}_{t}=\hat{H}^{(0)}+\hat{H}^{(1)}_{-1}e^{-i\omega t}+\hat{H}^{(1)}_{1}e^{i\omega t}$, by expanding the expectation value $\langle\hat{\mathcal{U}}_{t,t_{0}}^{{\dagger}}\hat{A}\hat{\mathcal{U}}_{t,t_{0}}\rangle$ in orders of $g$ and picking out terms oscillating at $e^{ij{\omega}t}$ where $j\in\mathds{Z}$, we find that the $\mathcal{O}(g)$ linear response of an observable $\hat{A}$ is given by $\delta^{(1)}\langle\hat{A}\rangle=\langle[\hat{F}^{(1)}_{t},\hat{A}]\rangle$, and the $\mathcal{O}(g^{2})$ response is given by $\delta^{(2)}\langle\hat{A}\rangle=\langle[\hat{F}^{(2)}_{t},\hat{A}]\rangle+\frac{1}{2}\langle[\hat{F}^{(1)}_{t},[\hat{F}^{(1)}_{t},\hat{A}]]\rangle$. This means we can directly compute these responses when $\hat{f}^{(n)}_{j}$ is found..
@@ D.3 The $\mathcal{O}(g^{3})$ third-lowest order
where we have defined $\beta_{4}\equiv\beta_{3}+\beta_{2}/24+\beta^{\prime}/6$, $\gamma_{4}\equiv\gamma_{3}+\gamma_{2}/24+\gamma^{\prime}/6$ and $\delta_{4}\equiv\delta_{3}+\delta_{2}/24+\delta^{\prime}/6$. In the non-interacting case $U=0$, these 3 parameters vanish, and then $\hat{H}^{\prime(4)}$ reduces exactly to the non-interacting dynamical localisation effect predicted before in Ref. [60]: The bandwidth renormalisation factor $\frac{1}{4}g^{4}/{\omega}^{4}$ in Eq. (65) matches the Taylor expansion of the Bessel function $J_{0}(2g/{\omega})=1-g^{2}/{\omega}^{2}+\frac{1}{4}g^{4}/{\omega}^{4}+\mathcal{O}(g^{6})$.
@@ E.2 The accuracy of FSWT Hamiltonian at low frequencies
$\begin{split}&\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j+1,\bar{s}}\mathcal{P}\\ &=\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j+1,\bar{s}}(1-\hat{n}_{j+1,s}\hat{n}_{j+1,\bar{s}})\mathcal{P}=0,\end{split}$ (66)
and similarly $\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j,\bar{s}}\hat{n}_{j+1,\bar{s}}\mathcal{P}=0$. Furthermore, since we consider the half-filling case, we have $\hat{n}_{j,\bar{s}}\to 1-\hat{n}_{j,s}$, which results in
$\begin{split}\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j,\bar{s}}\mathcal{P}&\to\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}(1-\hat{n}_{j,s})\mathcal{P}\\ &=\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\mathcal{P}.\end{split}$ (67)

### v3

@@ VI Acknowledgement
The authors would like to thank André Eckardt, Thomas Pohl, Francesco Petiziol, Alexander Schnell, and Li You for their helpful discussions on the spurious Floquet-mediated interactions and the relation with van Vleck theory. FPMMC, DJ and FS acknowledge support from the Cluster of Excellence ’Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994. DJ acknowledges support from the Hamburg Quantum Computing Initiative (HQIC) project EFRE. The project is co-financed by ERDF of the European Union and by “Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB)”.
@@ A.1 Link to the retarded Green function
where $\theta_{t}$ is the step function at $t=0$, $\big{(}\hat{H}_{1}^{(1)}\big{)}^{H}_{t}=e^{i\hat{H}^{(0)}t}\hat{H}_{1}^{(1)}e^{-i\hat{H}^{(0)}t}$ represents the Heisenberg operator of $\hat{H}_{1}^{(1)}$. Hence, $\langle[\hat{f}^{(1)}_{1},\hat{A}]\rangle$ is a frequency-domain retarded Green function of the undriven system $\hat{H}^{(0)}$. This direct link between $\hat{f}^{(n)}_{j}$ and the retarded Green function suggests that we can obtain the linear and non-linear responses of the system from the solutions of the Sylvester equations   [^6] For example, in a monochromatically driven system described by $\hat{H}_{t}=\hat{H}^{(0)}+\hat{H}^{(1)}_{-1}e^{-i\omega t}+\hat{H}^{(1)}_{1}e^{i\omega t}$, by expanding the expectation value $\langle\hat{\mathcal{U}}_{t,t_{0}}^{{\dagger}}\hat{A}\hat{\mathcal{U}}_{t,t_{0}}\rangle$ in orders of $g$ and picking out terms oscillating at $e^{ij{\omega}t}$ where $j\in\mathds{Z}$, we find that the $\mathcal{O}(g)$ linear response of an observable $\hat{A}$ is given by $\delta^{(1)}\langle\hat{A}\rangle=\langle[\hat{F}^{(1)}_{t},\hat{A}]\rangle$, and the $\mathcal{O}(g^{2})$ response is given by $\delta^{(2)}\langle\hat{A}\rangle=\langle[\hat{F}^{(2)}_{t},\hat{A}]\rangle+\frac{1}{2}\langle[\hat{F}^{(1)}_{t},[\hat{F}^{(1)}_{t},\hat{A}]]\rangle$. This means we can directly compute these responses when $\hat{f}^{(n)}_{j}$ is found.. Practically, to evaluate Eq. (37) for strongly correlated systems, we need to obtain the thermal density operator numerically, which we leave for future work.
@@ D.3 The $\mathcal{O}(g^{3})$ third-lowest order
where we have defined $\beta_{4}\equiv\beta_{3}+\beta_{2}/24+\beta^{\prime}/6$, $\gamma_{4}\equiv\gamma_{3}+\gamma_{2}/24+\gamma^{\prime}/6$ and $\delta_{4}\equiv\delta_{3}+\delta_{2}/24+\delta^{\prime}/6$. In the non-interacting case $U=0$, these 3 parameters vanish, and then $\hat{H}^{\prime(4)}$ reduces exactly to the non-interacting dynamical localisation effect predicted before in Ref. [64]: The bandwidth renormalisation factor $\frac{1}{4}g^{4}/{\omega}^{4}$ in Eq. (65) matches the Taylor expansion of the Bessel function $\mathcal{J}_{0}(2g/{\omega})=1-g^{2}/{\omega}^{2}+\frac{1}{4}g^{4}/{\omega}^{4}+\mathcal{O}(g^{6})$.
D.4 FSWT in the strong driving frame
To apply FSWT to the strongly driven Hubbard model (i.e., $g\gg{\omega}$ ) it is favourable to change into a strong driving frame as, otherwise, we would require to solve the Sylvester equations to higher orders of $g$. The change into a strong driving frame is a direct route to observe the Bessel function renormalization in the FSWT Hamiltonian, as we explain below. To change frames, we apply the Peierls transform [92, 94, 95]
$\hat{U}^{\text{r}}_{t}=\exp\left(i\frac{2g}{{\omega}}\sin({\omega}t)\sum_{j=1}^{L}j\sum_{s}\hat{n}_{j,s}\right)$ (66)
to the lab-frame Hamiltonian in Eq. (18). The transformed Hamiltonian acquires a Peierls phase coupling, replacing the driving strength $g$ with $J$. It reads
$\begin{split}\hat{H}_{P,t}=\hat{H}_{P}^{(0)}+\sum_{j=-\infty}^{\infty}\hat{H}^{(1)}_{P,j}e^{ij\omega t},\end{split}$ (67)
where the superscript now represents the orders of hopping $J$. In Eq. (67), the undriven part no longer contains hopping, as it is given by
$\hat{H}_{P}^{(0)}=U\sum_{{\bf j}}\hat{n}_{{\bf j},\uparrow}\hat{n}_{{\bf j},\downarrow},$ (68)
and the driving term oscillating at $e^{ij\omega t}$ reads
$\hat{H}^{(1)}_{P,j}=-J~{}\mathcal{J}_{j}(2g/\omega)\sum\limits_{s}\sum\limits_{j=1}^{L-1}(\hat{c}_{j+1,s}^{{\dagger}}\hat{c}_{j,s}+\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}),$ (69)
where $\mathcal{J}_{j}$ is the $j$th kind Bessel function. In this rotating frame, the driving term has a static component $\hat{H}^{(1)}_{P,0}$ which accounts for the well known $\mathcal{J}_{0}(2g/{\omega})$ bandwidth renormalisation. The $\mathcal{O}(J)$ lowest order Sylvester equation for the Hamiltonian in Eq. (67) is, for all $j\neq 0$,
$\hat{H}^{(1)}_{P,j}+[\hat{f}^{(1)}_{j},\hat{H}_{P}^{(0)}]-\omega\hat{f}^{(1)}_{j}=0.$ (70)
We thus see that, solving Eq. (70) is equivalent to solving Eq. (25b), which results in the parameters $\beta^{\prime}$, $\gamma^{\prime}$ and $\delta^{\prime}$. The corresponding FSWT Hamiltonian in this strong driving frame will be explored in future work.
@@ E.2 The accuracy of FSWT Hamiltonian at low frequencies
Besides, to benchmark the accuracy of FSWT in the in-gap driving regime where ${\omega}<U$, in Fig. 7 we simulate the return rate for $U=28J$ with other parameters matching Fig. 3(b). In this in-gap regime, our FSWT result (29) again outperforms the lowest-order HFE result, while higher-order lab-frame HFE corrections lead to divergence, rendering them ineffective.
Finally, we benchmark the accuracy of FSWT in predicting the response of physically observable quantities. In Fig. 8, we plot the driven dynamics of the nearest-neighbour doublon-doublon correlation at the centre of the chain, using the same parameters as Fig. 3(b). We again start the simulation from the product CDW state $|\psi_{CDW}\rangle$, thus the nearest-neighbour doublon correlation is initially zero. Compared to the exact TEBD simulation, the FSWT stroboscopic Hamiltonian (29) predicts the correlation dynamics with significantly improved accuracy than HFE. This improvement arises solely from the driving-induced $\mathcal{O}(J)$ and $\mathcal{O}(J^{2})$ correlated hoppings in Eq. (29).
Figure 7: The return rate Eq. (30) in the in-gap driving regime, with $L=10$, $\omega=16J$, $U=28J$ and $g={\omega}/4$.
Figure 8: The doublon-doublon correlation $\langle\hat{D}_{j}\hat{D}_{j+1}\rangle$ for the parameters considered in Fig. 3(b), with $L=10$, $\omega=16J$, $U=8J$ and $g={\omega}/4$. Here $\hat{D}_{j}=\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}$ is the doublon density operator, and we take $j=L/2$ to observe the doublon-doublon correlation at the centre of the chain.
E.3 The accuracy of FSWT for other initial states
The FSWT result remains valid in predicting the dynamics starting from other initial states. For example, the return rate starting from a far-from-equilibrium product state with 2/3-filling is shown in Fig. 9. Again, the FSWT stroboscopic dynamics outperforms the HFE result throughout the dynamics.
Figure 9: The return rate Eq. (30) starting from a non-half-filling state. The parameters match those used in Fig. 3(c), with $L=6$, $\omega=10J$, $U=3J$ and $g={\omega}/4$.
@@ E.3 The accuracy of FSWT for other initial states
$\begin{split}&\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j+1,\bar{s}}\mathcal{P}=\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j+1,\bar{s}}(1-\hat{n}_{j+1,s}\hat{n}_{j+1,\bar{s}})\mathcal{P}=0,\end{split}$ (71)
and similarly $\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j,\bar{s}}\hat{n}_{j+1,\bar{s}}\mathcal{P}=0$. Furthermore, since we consider the half-filling case, we have $\hat{n}_{j,\bar{s}}\to 1-\hat{n}_{j,s}$, which gives
$\begin{split}\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\hat{n}_{j,\bar{s}}\mathcal{P}&\to\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}(1-\hat{n}_{j,s})\mathcal{P}=\hat{c}_{j,s}^{{\dagger}}\hat{c}_{j+1,s}\mathcal{P}.\end{split}$ (72)